OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?" 1 4 −10 −6

OpenStudy (nurali):

y = (x − 3)(4x + 2) y = 4x² + 2x − 12x − 6 y = 4x² − 10x − 6 y = ax² + bx + c ⇒ b = −10